OpenStudy (lilai3):

@whpalmer4 -2 +2x^2 Factor the polynomial. Please help for a medal....or an oreo ;) <3

OpenStudy (whpalmer4):

\[-2+2x^2\]Is that it?

OpenStudy (lilai3):

I didn't solve this one yet. Yes, that's the problem! :)

OpenStudy (whpalmer4):

Okay. See any common factors?

OpenStudy (lilai3):

Uh...um...2?

OpenStudy (whpalmer4):

Any others?

OpenStudy (lilai3):

-2?

OpenStudy (whpalmer4):

that's a better one. What do you have after you factor that out?

OpenStudy (lilai3):

Uh... -2 (1-2x^2) Is that right?

OpenStudy (whpalmer4):

not quite. \[-2(1-2x^2) = -2*1 -2*(-2)x^2 = -2+4x^2\]which is not what you started with.

OpenStudy (lilai3):

What? I do not understand what you just did...please explain?

OpenStudy (whpalmer4):

I just multiplied out what you factored. I should get the same thing we started with if neither of us made any mistakes.

OpenStudy (whpalmer4):

I didn't, but you did :-)

OpenStudy (lilai3):

Wait, what? Okay, can we just start over. The common factor is -2. Then...?

OpenStudy (lilai3):

Because it's really confusing

OpenStudy (whpalmer4):

Factor -2 out of \(-2+2x^2\). In other words, what do you have to multiply by -2 to get \(-2+2x^2\)?

OpenStudy (lilai3):

-2 (1+x^2)

OpenStudy (whpalmer4):

Let's do it term by term: what is \[\frac{-2}{-2}=\]

OpenStudy (lilai3):

\[-2 (1+x ^{2})\]

OpenStudy (whpalmer4):

No, answer my question, please...what is \[\frac{-2}{-2}\]

OpenStudy (lilai3):

-2/-2 equals to 1.

OpenStudy (whpalmer4):

Okay, so the first term is 1 What is \[\frac{2x^2}{-2}=\]

OpenStudy (lilai3):

-1x^2

OpenStudy (whpalmer4):

Right! so the second term is \(-x^2\), giving us\[-2+x^2 = -2(1-x^2)\]If we multiply that out, we'll get what we started with: \[-2(1-x^2) = -2*1 -2*(-x^2) = -2 +2x^2\checkmark\]

OpenStudy (whpalmer4):

Sorry, that first line should have been \[-2 + 2x^2 = -2(1-x^2)\]

OpenStudy (lilai3):

So I am right. Yay!(:

OpenStudy (whpalmer4):

No, you had \[-2+2x^2 = -2(1+x^2)\]and that is not right.

OpenStudy (lilai3):

no, i put a minus sign

OpenStudy (lilai3):

oh wait never mind my mistake...i meant to though

OpenStudy (whpalmer4):

Okay :-) So, are we done factoring?

OpenStudy (lilai3):

uh no

OpenStudy (whpalmer4):

Good. what's the next step?

OpenStudy (whpalmer4):

Hint: Don't pull out another problem :-)

OpenStudy (lilai3):

What? -2 (1+x) (1-x) ? Is that right?

OpenStudy (whpalmer4):

Well, let's check: \[-2(1+x)(1-x) = -2(1*1 -1*x + 1*x -x*x ) = -2(1-x^2) = -2+2x^2\checkmark\]

OpenStudy (whpalmer4):

@#$@# OS web design, gives you more space when typing than will actually show up on the screen when posted :-(

OpenStudy (lilai3):

LOL

OpenStudy (whpalmer4):

Yes, that is correct.

OpenStudy (lilai3):

Thanks. Next problem? I'll tag you. You have been such a great help so far. Thank you so much!

OpenStudy (whpalmer4):

you had a bit of trouble factoring out the -2, but you made up for it by spotting the backwards difference of squares, I was very pleased :-)

OpenStudy (lilai3):

My grandpa helped me ...XP

OpenStudy (lilai3):

LOL thank you though